Find the sum of two vectors A=5m<60°and B=6m<130°


1
Expert's answer
2021-12-23T16:43:45-0500

Find the components:

"A_x=5\cos{60°}=2.5\text{ m};\\
A_y=5\sin{60°}=4.33\text{ m};\\
B_x=6\cos{130°}=-3.9\text{ m};\\
B_y=6\sin{130°}=4.6\text{ m};\\
R_x=A_x+B_x,\\
R_y=A_y+B_y,\\
R=✓(R_x^2+R_y^2)=9\text{ m},\\
\theta=\arctan(R_y/R_x)=-8.9°."